OpenStudy (anonymous):

Write the standard equation of the circle below. Answer (x + 1)2 + (y - 2)2 = 1 (x - 1)2 + (y - 2)2 = 1 (x + 1)2 + (y + 2)2 = 1 (x - 1)2 + (y - 2)2 = 1

OpenStudy (imstuck):

Do you have an image of the circle to post?

OpenStudy (anonymous):



OpenStudy (anonymous):

sorry its bad its at -1,2

OpenStudy (imstuck):

I'm sorry, but that is kind of a bad drawing...is the center at (-1, 2)? Not that my drawing is any better... ; )

OpenStudy (anonymous):

yea it is

OpenStudy (imstuck):

The standard equation for a circle is\[(x-h)^{2}+(y-k)^{2}=r ^{2}\]Are you familiar with that?

OpenStudy (anonymous):

somewhat

OpenStudy (imstuck):

The h and the k indicate that the center is something other than the origin which is (0, 0).

OpenStudy (anonymous):

ok

OpenStudy (imstuck):

Your center is at (-1, 2), so you know right away some of what you need to begin filling in the equation, which is to fill in -1 for h and 2 for k). But keep in mind, the signs are opposite from what your coordinate reflects. Like this:

OpenStudy (imstuck):

\[(x+1)^{2}+(y-2)^{2}=r ^{2}\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

yea i get that part now

OpenStudy (imstuck):

This is because the standard form must be written in (x - h) and (y - k). So if your center x coordinate is -1, it fits into the formula like this (x - (-1)) which equals (x + 1). Likewise for the y coordinate. It fits into the formula like this: (y - (+2)) which equals (y - 2). See that?

OpenStudy (imstuck):

Now we just need to determine the radius. Let me get a better look at your circle you drew...

OpenStudy (anonymous):

but its bad

OpenStudy (imstuck):

Do we have a point on the circle? I have to know a point on the circle in order to determine the radius.

OpenStudy (anonymous):

and yea i understand the (x+1) + (y-1)=r

OpenStudy (anonymous):

it doesn't say

OpenStudy (imstuck):

Don't forget those are squared terms. And if it doesn't say on your work, I need you to tell me something that only you can determine from the picture. Do you actually have a drawing or an image that you took the circle from?

OpenStudy (anonymous):

um i doont have one

OpenStudy (imstuck):

Then what do you have? A picture? A list of specifics about the circle? Anything other than the center?

OpenStudy (imstuck):

Here is what I am talking about. Let me post my own pic as an example:

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

oh its not asking for radius

OpenStudy (imstuck):

You don't need to complete the circle's equation, then?

OpenStudy (anonymous):

no (x + 1)2 + (y - 2)2 = 1 (x - 1)2 + (y - 2)2 = 1 (x + 1)2 + (y + 2)2 = 1 (x - 1)2 + (y - 2)2 = 1

OpenStudy (imstuck):

I get it! Which one is it from that list, then? We determined the center, and obviously they gave you the radius, which is 1, so which of those is the correct one?

OpenStudy (anonymous):

the x=1 y-2 ?

OpenStudy (anonymous):

so a?

OpenStudy (imstuck):

It's the first choice in your list. Just remember how to find the center AND then how to write it into the equation and you'll be fine!

OpenStudy (imstuck):

Good job!

OpenStudy (imstuck):

And thank you for the medal!

OpenStudy (anonymous):

thx i medaled and fanned u u wl be hearing from me again because ur smart

OpenStudy (imstuck):

Actually, I teach this every day...I'm not that smart...just experienced! And thank you!

OpenStudy (anonymous):

well ur good at it